4. The number of spectators that turn up to support their local cricket team is known to have a Poisson distribution with mea(b) Determine the expected profit (Write your answer in the box provided and show any working below.)(c) Determine the variance of the profit. (Write your answer in the box provided and show any working below.)
Answer #1

Let X represent the number of people who turn up. Therefore the distribution for X here is given as:

XPoisson(10)

Now the profit is computed as:

= Total Revenue - Total Cost

= 3X - 16

because 16 is the total cost that is the gatekeeper's fees.

a) Probability that there is a profit is computed as:

P( 3X > 16)

= P(X > 16/3)

= 1 - P(X <= 5)

This is computed using the poisson probability function as:

102 103 104 105 -10 -10 -10 e 2 -10 e 3! -10 e -10 e 1 -6 10e - 4! 5!

0.93291

Therefore 0.93291 is the required probability here.

b) The expected value of the profit is computed as the expected value of the random variable P as:

E(P) = E(3X - 16) = 3E(P) - 16 = 3*10 - 16 = 14

Therefore $14 is the expected profit here.

c) The variance of the profit here is computed as:

Var(P) = Var(3X - 16) = 32 Var(X) = 9*10 = 90

Note that the variance for a poisson random variable is same as the mean of that variable.